# Scaling Intelligence: Designing Data Centers for Next-Gen Language Models

### v1

@@ 3.11. Performance Benefit of FullFlat Network
Our studies show FullFlat networks offer superior performance and allows better utilization of increased resouces. This setup allows unconventional strategies, such as using TP values going beyond a node boundary (suggested by popular wisedom), typically avoided in Two-Tier networks, especially with smaller HBD sizes. FullFlat with 1.6 TB/s bandwidth at 80% efficiency improves performance by 30-60% over Two-Tier networks with 200 GB/s scale-out bandwidth, enhancing MFU (see Figure 8) and system utilization to 70% or more. Since FullFlat topology is less sensitive to missing optimizations, like comp/comm overlaps, users benefit from reduced programming effort, as the top 5,000 configurations show only a 5% performance gap (see Figure 1). Programmers also need to worry less about achieving perfect comp/comm overlaps.
A Full Flat network with advanced CPO offers improved reliability, availability, and serviceability. Fully optical interconnects are expected to experience fewer link flaps and would require fewer switches resulting in reduced power and costs (Patel et al., 2025). A Full Flat network improves resiliency, and robustness, and requires less redundancy tax since any GPU can replace any faulty one anywhere in the network. With equal HBD/LBD bandwidth and a high-radix, low-diameter, low-latency network, if a GPU fails in an HBD, it can be easily replaced with a spare from anywhere in the cluster, rather than replacing the entire HBD domain, improving TCO in the data center. While detailed TCO/power modeling is outside this paper’s scope, at a high level, CPO optics deliver $\approx~{}30-50\%)lower$/GB/s than discrete optics, Co-packaged optics reduce per-port power by 30–50% (Chang, 2024a, b, c). Thus, FullFlat can reduce data center TCO and energy usage by up to 20% or even more with innovations.

### v2

@@ 3.11. Performance Benefit of FullFlat Network
Our studies show FullFlat networks offer superior performance and allows better utilization of increased resources. This setup allows unconventional strategies, such as using TP values going beyond a node boundary (suggested by popular wisdom), typically avoided in Two-Tier networks, especially with smaller HBD sizes. FullFlat with 1.6 TB/s bandwidth at 80% efficiency improves performance by 30-60% over Two-Tier networks with 200 GB/s scale-out bandwidth, enhancing MFU (see Figure 8) and system utilization to 70% or more. Since FullFlat topology is less sensitive to missing optimizations, like comp/comm overlaps, users benefit from reduced programming effort, as the top 5,000 configurations show only a 5% performance gap (see Figure 1). Programmers also need to worry less about achieving perfect comp/comm overlaps.
A Full Flat network with advanced CPO offers improved reliability, availability, and serviceability. Fully optical interconnects are expected to experience fewer link flaps and would require fewer switches resulting in reduced power and costs (Patel et al., 2025). A Full Flat network improves resiliency, and robustness, and requires less redundancy tax since any GPU can replace any faulty one anywhere in the network. With equal HBD/LBD bandwidth and a high-radix, low-diameter, low-latency network, if a GPU fails in an HBD, it can be easily replaced with a spare from anywhere in the cluster, rather than replacing the entire HBD domain, improving TCO in the data center. While detailed TCO/power modeling is outside this paper’s scope, at a high level, CPO optics deliver $\approx~30-50\%$ lower GB/s than discrete optics, Co-packaged optics reduce per-port power by 30–50% (Chang, 2024a, b, c). Thus, FullFlat can reduce data center TCO and energy usage by up to 20% or even more with innovations.